OpenStudy (anonymous):

find the acce. of d moon ith respect 2 d earth 4rm d following data: distance b/w d earth & moon =3.85*10^5 km & the time taken by moon to complete one revolution around the earth =27.3 days.

OpenStudy (anonymous):

@apoorvk @FoolAroundMath @timo86m @lgbasallote plz hlp me out.

OpenStudy (anonymous):

hey @apoorvk hlp me out

OpenStudy (apoorvk):

The distance of the moon is the radius, and you have the time period as well, so you can find out the velocity of rotation. Now, a = (v^2)/r Use this to find out the acceleration!

OpenStudy (vincent-lyon.fr):

Please, use full-text when you type questions.

OpenStudy (anonymous):

you have radius r acceleration=angular acceleration*r and angular acceleration =2pir/t

OpenStudy (anonymous):

oops a=w^2*r where w=2.pi.r/t

OpenStudy (anonymous):

i hav done this

OpenStudy (anonymous):

@theyatin

OpenStudy (anonymous):

thats good :)

OpenStudy (anonymous):

after that i'm nt getting answer

OpenStudy (anonymous):

@ParthKohli

OpenStudy (apoorvk):

Did you take care of the units? 27.3 *days*